OpenStudy (anonymous):

Deb and Sam go to the mall, each have $200 to spend. Deb spent $150 and Sam spent $160. What percent of their total money did they spend? A. 57.5% B. 75.7% C. Cannot calculate the percentage with the information provided. D. 77.5%

OpenStudy (anonymous):

D

OpenStudy (anonymous):

% = part/whole part = 150 + 160 = 310 whole = 200 + 200 = 400 310/400 * 100 that is how you calculate, but Bookreader is correct